On December 31, Pacifica, Inc., acquired 100 percent of the voting stock of Seguros Company. Pacifica will maintain Seguros as a wholly owned subsidiary with its own legal and accounting identity. The consideration transferred to the owner of Seguros included 56,500 newly issued Pacifica common shares ($20 market value, $5 par value) and an agreement to pay an additional $130,000 cash if Seguros meets certain project completion goals by December 31 of the following year. Pacifica estimates a 50 percent probability that Seguros will be successful in meeting these goals and uses a 4 percent discount rate to represent the time value of money Immediately prior to the acquisition, the following data for both firms were available Seguros Seguros Pacifica Book Values Fair Values $ (2,110,000) Revenues Expenses 1 , 477, 000 Net income $ (633,000) Retained earnings, 1/1 Net income Dividends declared $ (1,026,0e) (633,000) 171,000 Retained earnings, 12/31 Cash Receivables and inventory Property, plant, and equipment Trademarks $(1,488,000) $ 162,000 154,000 $154,000 73,500 662,50e 293,000e 93,000 487,000 248,00e 2,959,00 $ 982,000 254,0080 2,190,0ee 353,00e Total assets Liabilities Common stock Additional paid-in capital Retained earnings $ (596,00) (258,000) (258,000) (400,000) (200,000) (7,000) (1,488,000) (454,008 $(2,959,000) (982,000) (475,000) Total liabilities and equities In addition, Pacifica assessed a research and development project under way at Seguros to have a fair value of $157,000. Although not yet recorded on its books, Pacifica paid legal fees of $24,600 in connection with the acquisition and $11,700 in stock issue costs a. Prepare Pacificas entries to account for the consideration transferred to the former owners of Seguros, the direct combination costs, and the stock issue and registration costs b.&c. Present a worksheet showing the postacquisition column of accounts for Pacifica and the consolidated balance sheet as of the acquisition date
Answer #1

Part A

Account titles and explanation

debit

credit

Investment in Seguros

1192500

Common Stock (56500 × $5)

282500

Additional Paid-In Capital (56500 × $15)

847500

Contingent performance obligation (130000*50%*0.961538)

62500

(to record the acquisition on Pacifica’s records)

Professional Services Expense

24600

Cash

24600

Additional Paid-In Capital

11700

Cash

11700

Part B & C

Pacifica

Seguros

Consolidation Entries

Consolidated Balance Sheet

Revenues

(2110000)

(2110000)

Expenses

1477000

1477000

Net income

(633000)

(633000)

Retained earnings, 1/1

(1026000)

(1026000)

Net income

(633000)

(633000)

Dividends declared

171000

171000

Retained earnings, 12/31

(1488000)

(1488000)

Cash

125700

154000

279700

Receivables and inventory

254000

93000

19500

327500

Property, plant and equipment

2190000

487000

175500

2852500

Investment in Seguros

1228800

1228800

0

Research and development asset

157000

157000

Goodwill

146800

146800

Trademarks

353000

248000

45000

646000

Total assets

2959000

982000

(4409500)

Liabilities

(596000)

(258000)

(854000)

Contingent performance obligation

(62500)

(62500)

Common stock

(682500)

(200000)

200000

(682500)

Additional paid-in capital

(1322500)

(70000)

70000

(1322500)

Retained earnings

(1488000)

(454000)

454000

(1488000)

Total liabilities and equities

(2959000)

(982000)

(4409500)

162000-24600-11700 = 125700

93000-73500 = 19500

662500-487000 = 175500

293000-248000 = 45000

400000+282500 = 682500

475000+847500 = 1322500